OpenStudy (anonymous):

some one plz help me i will give medal Given the function f:x=5-2x, evaluate each of the following. (1) f(3)+f(-3)

OpenStudy (mayankdevnani):

\[\large \bf f(x)=5-2x\] then plug x=3, we get \[\large \bf f(3)=5-2(3)=5-6=-1\] then,plug x=-3,we get \[\large \bf f(-3)=5-2(-3)=5+6=11\]

OpenStudy (mayankdevnani):

and then solve !!! @zanubiah

OpenStudy (anonymous):

ya i m here i was checking ur profile

OpenStudy (anonymous):

ok! thank u so much ya i got my answer

OpenStudy (mayankdevnani):

welcome :)